At the fair, you buy 3 sausage sandwiches and a milkshake and it costs you $8.25. Your friend buys 1 sausage sandwich and 2 milkshakes and her total is $5.25. What is the cost of one sandwich and one milkshake?

A. Write a system of equations for this situation.

B. Solve your system using either elimination or substitution. Make sure to show your work for full credit.

Answer by amarjeeth123(570)

Let the cost of one sausage sandwich be x.
Let the cost of one milkshake be y
We get,
3x+y=8.25......equation 1
x+2y=5.25......equation 2
Multiplying equation 1 by 2 we get,
6x+2y=16.5.....equation 3
Subtracting equation 2 from 3 we get,
5x=11.25
x=2.25
y=8.25-6.75=1.5
The cost of one sausage sandwich is $2.25 and the cost of one milkshake is $1.5.